OpenStudy (anonymous):

how do you use the half life equation while trying to find the begining ammount?

OpenStudy (anonymous):

u may have to use the natural logarithm / ln

OpenStudy (anonymous):

ok but u need to supply with info though

OpenStudy (anonymous):

sorry heading to bed now so won't be able to help u :( sorry abt that

OpenStudy (anonymous):

Let's say the half life equation is \[N(t) = N_0 (\frac{1}{2}) ^\frac{t}{Th}.\] Then the inital amount N_0 is solved for, this way \[N_0 = N(t) \frac{1}{2}^{-\frac{t}{Th}}.\] If you have time:t, current amount:N(t), and half life Th, N_0 can be calculated.

OpenStudy (dumbcow):

Sometimes the equation may also be represented as P(t) = P_0*e^-kt P_0 = P(t)*e^kt, where k can be found if given the half life